OpenStudy (anonymous):

3x^2-3x-1>-1 please solve using quadratic inequalities

OpenStudy (anonymous):

start with \[3x^2-3x>0\] and factor as \[3x(x-1)>0\] this one is zero at 0 and 1, and since it is a parabola that opens up it is positive outside the zeros, namely if \(x<0\) or if \(x>1\)

OpenStudy (anonymous):

i really appriciate it. you have no idea:)

OpenStudy (anonymous):

it is clear (more or less) how to do it? you have to have a zero on one side of the inequality then you need to find the actual zeros then think about whether you want positive or negative

OpenStudy (anonymous):

we can work through another one if you have one more to do

OpenStudy (anonymous):

its clear! i just want to work one more because the sign > has a line underneath it

OpenStudy (anonymous):

-x^2+x+2>0

OpenStudy (anonymous):

then so will your answer

OpenStudy (anonymous):

i hate working with negative leading coefficients, so if this is \[-x^2+x+2\geq 0\] then i would change the sign and the inequality and start with \[x^2-x-2\leq 0\]

OpenStudy (anonymous):

again we factor, this time as \[(x+1)(x-2)\] so the zeros are at \(x=-1\) and at \(x=2\)

OpenStudy (anonymous):

we are looking for where it is less than or equal to zero, and again this is a parabola that faces up, so it is negative between the zeros i.e. if \[-1\leq x\leq 2\]

OpenStudy (anonymous):

it is much more convenient to work with a positive leading coefficient, rather than a negative one that is why i changed \[-x^2+x+2\geq 0\] to \[x^2-x-2\leq 0\]

OpenStudy (anonymous):

wow i see that makes sense!!

OpenStudy (anonymous):

i am so happy that i found this site

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

you are a great person for helping others :) thank you this really makes sense!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if you dont mind helping me with one more?

OpenStudy (anonymous):

4x^2+4x>3

OpenStudy (anonymous):

the > has a line underneath it also